what are the missing numbers 3/4 ? 27/64 81/256 ? 729/4096
Answer by dhartman0220(8)

the series is increasing exponentially.... (3/4)^x
x=1, x=2, x=3, x=4.....x
(3/4)^1; (3/4)^2; 3/4^3; (3/4)^4; (3/4)^5; (3/4)^6
(3/4)^2=9/16
(3/4)^5=243/1024